OpenStudy (goformit100):

Solve: 3(x + 4) greater than or equal to x - 2 ? x greater than or equal to -7 x greater than or equal to 7 x less than or greater to -7 x greater than or equal to -3

OpenStudy (goformit100):

@stgreen

OpenStudy (goformit100):

@zzr0ck3r

OpenStudy (anonymous):

first one

OpenStudy (zzr0ck3r):

3(x+4)>=x-2 3x+12>=x-2 2x+12>=-2 2x>=-14 x>=-7

OpenStudy (goformit100):

Thanks